4) Suppose the demand and supply curves for eggs are given by the following equations:
Qd = 100-20P
Qs = 10 + 40P
a) Fill up the following:
Price Qd Qs
.50 _____ _____
1.00 _____ _____
1.50 _____ _____
2.00 _____ _____
2.50 _____ _____
b)Use the information in the table to find the equilibrium price and quantity
c) Graph the demand supply curves and identify the equilibrium price and quantity
1
Expert's answer
2021-09-16T16:56:03-0400

When the price ;

"0.5"

"Qd = 100-20(0.5) = 90"

"QS = 10+40(0.5) = 30"


"1"

"Qd = 100-20(1) = 80"

"QS = 10 + 40(1) = 50"


"1.5"

"Qd = 100-20(1.5) = 70"

"QS = 10+40(1.5) = 70"


"2"

"Qd = 100-20(2) = 60"

"QS = 10 + 40(2) = 90"


"2.5"

"Qd = 100-20(2.5) = 50"

"QS = 10 + 40(2.5) = 110"


Equilibrium price

"100-20p = 10+40p"


"100-10 = 40p +20p"


"90 = 60P"

"P = \frac{90}{60} = 1.5"

Equilibrium price = 1.5


Equilibrium quantity

"Qd = 100-20p"


"= 100-20(1.5) = 70"

"Equilibrium quantity = 70"